# RelCon: Relative Contrastive Learning for a Motion Foundation Model for Wearable Data

### v4

@@ Abstract
We present RelCon, a novel self-supervised Relative Contrastive learning approach that uses a learnable distance measure in combination with a softened contrastive loss for training an motion foundation model from wearable sensors. The learnable distance measure captures motif similarity and domain-specific semantic information such as rotation invariance. The learned distance provides a measurement of semantic similarity between a pair of accelerometer time-series segments, which is used to measure the distance between an anchor and various other sampled candidate segments. The self-supervised model is trained on 1 billion segments from 87,376 participants from a large wearables dataset. The model achieves strong performance across multiple downstream tasks, encompassing both classification and regression. To our knowledge, we are the first to show the generalizability of a self-supervised learning model with motion data from wearables across distinct evaluation tasks.
@@ 1 Introduction
Advances in self-supervised learning (SSL) combined with the availability of large-scale datasets have resulted in a proliferation of foundation models (FMs) in computer vision (Oquab et al., 2023), NLP (OpenAI et al., 2023), and speech understanding (Yang et al., 2024). These models provide powerful, general-purpose representations for a particular domain of data, and support generalization to a broad set of downstream tasks without the need for finetuning. For example, the image representation contained in the DINOv2 (Oquab et al., 2023) model was trained in an entirely self-supervised way and achieves state-of-the-art performance on multiple dense image prediction tasks such as depth estimation and semantic segmentation, by decoding a frozen base representation with task-specific heads. In contrast to these advances, the times-series have not yet benefited from the foundation model approach, with a few exceptions (Abbaspourazad et al., 2024; Das et al., 2023). This is particularly unfortunate for problems in mobile health (mHealth) signal analysis, which encompasses data modalities such as accelerometry, PPG, and ECG (Rehg et al., 2017), as the collection of mHealth data from participants can be time-consuming and expensive. However, recent advances in self-supervised learning for mHealth signals (Abbaspourazad et al., 2024; Yuan et al., 2024; Xu et al., 2024) have shown promising performance, raising the question of whether it is now feasible to train foundation models for mHealth signals.
@@
In this paper, we demonstrate, for the first time, the feasibility of adopting a foundation model approach for the analysis of accelerometry data across tasks. Accelerometry is an important mHealth signal modality that is used in human activity recognition (HAR) (Haresamudram et al., 2022), physical health status assessment (Xu et al., 2022), energy expenditure estimation (Stutz et al., 2024), and gait assessment (Apple, 2021), among many other tasks. We use a novel method for self-supervised learning combined with pretraining on a large-scale accelerometry dataset. We show that this approach yields an effective representation for accelerometry data which is useful for multiple downstream tasks, including HAR and the estimation of gait metrics such as stride velocity and double support time. Crucially, we obtain state-of-the-art performance on these tasks without finetuning and also exceed the performance of fully-supervised models trained on smaller datasets.
The key to our approach is a novel method for contrastive representation learning which significantly extends the recent REBAR (Xu et al., 2024) architecture. Contrastive learning relies on the construction of positive and negative pairs that instruct the model to learn key differences and similarities between selected time-series segments. Prior works have explored a variety of methods for constructing positive and negative pairs. One straightforward approach is to select segments based on temporal proximity (Abbaspourazad et al., 2024; Kiyasseh et al., 2021; Jeong et al., 2023a). Unfortunately, this heuristic can fail when segments that are far apart in time have a similar structure, as in the case for repetitive or periodic movements. Another popular approach is to construct positive pairs via data augmentation (Haresamudram et al., 2022; Matton et al., 2023; Gopal et al., 2021). This method, exemplified by SimCLR (Chen et al., 2020), has been shown to be extremely effective for computer vision tasks, where the rich invariant structure stemming from the imaging process provides many sources for augmentations. Unfortunately, it has proven to be challenging to define common augmentations for general time-series data that are relevant across diverse downstream tasks and encode meaningful invariances, without inadvertently altering important signal properties.
Our approach exploits the observation that many time-series of interest are typically composed of motifs, short distinctive temporal shapes that may approximately repeat themselves over time. For example, the upward swing of the arm, captured by a wristworn accelerometer during a bout of walking, will repeat at regular intervals depending on the cadence. Prior work has exploited this observation to learn a distance measure that captures motif similarity (Xu et al., 2024), and used this measure to identify positive and negative pseudo-labels. Since this approach forms pairs by matching based on signal shape, it’s more likely to identify semantically meaningful pairs than approaches based on a heuristic like temporal distance, and since it can exploit natural variations in signal shapes across an extended time-series recording, it can identify within-person temporal dynamics, such as fatigue over time. Then, by introducing accelerometry-specific augmentations, we can improve the prior distance measure to learn to compare motifs, invariant to sensors positions or orientation. Another limitation of Xu et al. (2024) is that the contrastive loss is hard in the sense that all negative samples are treated equally in the loss computation. Accounting for the degree of similarity between the positive anchor sample and negative samples, via a soft loss, could imbue the embedding space with greater semantic meaning - for instance, having clusters of samples retain relative similarities like having ‘walking’ samples closer to ‘running’ than ‘yoga’ in an activity classification task. We introduce a novel RelCon method that accomplishes this goal and show that it significantly improves upon the performance of REBAR.
Figure 1: Comparisons between various SOTA SSL approaches for Accelerometry data. Each sequence color represents a different user’s time-series. SimCLR resists false positives by creating positive pairs through augmentations, preserving semantics, and models between-user interactions by drawing negatives from different users. RelCon’s loss function captures relative positions of the within- and between-user candidates with its learnable distance function, offering better resistance to false positives and negatives compared to binary comparison methods like REBAR.
@@
Our RelCon-trained motion foundation model achieves consistently strong performance across 6 different downstream datasets across 4 different types of downstream tasks for a wide set of metrics, outperforming current SOTA accelerometry SSL approaches. We are the first to show the generalizability of a self-supervised foundation model with motion data from wearables across each of these distinct evaluation tasks, ranging from activity classification to gait metric regression.
@@ 2 Related Work
Time-series FM: We define foundation models to be representations that are pre-trained on broad-scale data and are capable of solving multiple diverse downstream tasks without fine-tuning (Bommasani et al., 2021) (e.g., each task uses frozen weights with supervised training of a light-weight prediction head). We believe we are the first to train a model for accelerometry data which meets these criteria. The closest related work is Yuan et al. (2024), which uses data from the UK Biobank to train an accelerometry representation. However, this work does not exhibit multi-task downstream performance, since they are all Human Activity Recogintion (HAR) variants, making it is less clear how well their representation captures the data domain. In contrast, we test our FM on multiple diverse tasks including HAR, workout classification, and gait analysis regression. Other notable FMs are Abbaspourazad et al. (2024); Das et al. (2023), but these do not model accelerometry.
Time-series SSL: Our RelCon architecture is a novel contrastive learning approach for time-series data. The closest related method is REBAR (Xu et al., 2024). We adopt REBAR’s motif-based approach to generating positive pairs but their model was non-specific to accelerometry and thus only used a simple exact-motif-matching mechanism, which would not be invariant to changes in sensor position and other invariances. Additionally, due to its ”hard” contrastive loss, it suffers from sensitivity to false positives and negatives because only one candidate is set to be positive and all other are set to be negative. Other prior works on SSL primarily adopt either data augmentations or masked auto-encoder approaches, including other prior works on HAR (Yuan et al., 2024; Haresamudram et al., 2022; 2024; Straczkiewicz et al., 2021). Additional works have addressed SSL for signals such as PPG and ECG, for applications including health condition predictions and sleep staging (Abbaspourazad et al., 2024; Song et al., 2024; Thapa et al., 2024; McKeen et al., 2024; Yuan et al., 2024; Kiyasseh et al., 2021; Diamant et al., 2022; Jeong et al., 2023b; Das et al., 2023).
For motion wearables, most SSL approaches have focused on human activity recognition. Haresamudram et al. (2022) benchmarked a number of SSL approaches on accelerometer data, showing generalizability across datasets and sensor positions. Models for other motion-based tasks – including walking speed prediction (Shrestha & Won, 2018; Soltani et al., 2021), gait classification (Slemenšek et al., 2023; Brand et al., 2024), and health monitoring (Takallou et al., 2024) have focused on physics-based models, supervised learning, or SSL training for a single task.
@@
Soft Label Learning Approaches: A feature of our RelCon approach is the use of soft labels to obtain more fine-grained characterizations of similarity. Most prior self-supervised contrastive methods for time-series have adopted the hard label approach, including the previous REBAR method (Xu et al., 2024) as well as (Yue et al., 2022), (Tonekaboni et al., 2021), (Zhang et al., 2022), and (Abbaspourazad et al., 2024). One exception is Lee et al. (2023), which proposed a soft contrastive learning approach for downstream classification and anomaly detection, using non-learnable time-series distance measures, including dynamic time warping (Müller, 2007).
@@ 3 Methodology
The RelCon methodology has two key components. The first includes several innovations to learn a better distance measure to capture accel semantic information in Section 3.1. The second component is a novel relative contrastive loss that encodes relative order relationships, presented in Section 3.2.
@@ 3.1 Learnable Distance Measure
where $\smash{\textbf{X}\in\mathbb{R}^{T\times D}}$ and $\smash{\textbf{x},\mu,\sigma\in\mathbb{R}^{D}}$ with $T$ as the time length and $D$=3 for our 3-axis accelerometry signals. The distance between an anchor sequence and a candidate sequence, $d(\textbf{X}_{\textrm{anc}},\textbf{X}_{\textrm{cand}})$, is defined as the reconstruction accuracy to generate the anchor from the candidate. The distance measure is strictly dependent on the motif similarities between the anchor and candidate that are captured in the dilated convolutions in $\smash{f_{\{q/k\}}}$ (Xu et al., 2024).
@@
During training, we define $\textbf{X}_{\textrm{cand}}\coloneqq Aug(\textbf{X}_{\textrm{anc}})$, making the candidate an augmented version of the original sequence, using augmentations from Accel SimCLR (Tang et al., 2020). This helps the model learn to reconstruct the original anchor from semantically similar but altered candidates and prevents the exact-match solution, such as recognizing running signals even with changes like an upside-down wearable device or increased noise from a loose fitting device.
@@
    Modify the reversible instance normalization (Kim et al., 2021) to normalize an anchor based upon the candidate, to preserve relative magnitude information.
The anchor sequence and final reconstruction output are normalized (in Eq. 4) and unnormalized (in Eq. 2) using candidate sequence statistics. When an anchor and candidate have drastically different magnitudes, it is more likely they represent different fundamental motions and should have worse reconstruction error. Reversible normalization helps preserve this effect.
@@ 3.2 Relative Contrastive Loss: All pairs are positive …
Normalized temperature cross entropy loss (NT-Xent is standard in contrastive learning (Eq. 5). It learns a class discriminative embedding space by pulling the anchor and positive instances closer together in the embedding space and pushing all negatives away from the anchor, as in:
@@
Traditionally, NT-Xent uses strict labels that define absolute positive and negative sets. However, we would like to use the learned distance measure to modify the loss function to capture relative ordering among the candidates. That is, if $d(\textbf{X}_{\textrm{anc}},\textbf{X}_{i})>d(\textbf{X}_{\textrm{anc}},\textbf{X}_{j})$ the loss learns to preserve that relative ordering in the embedding space.
To do this, we redefine $\mathcal{S}_{\textrm{neg}}\coloneqq f_{\textrm{neg}}$ in Eq. 6. Now the set of negatives is replaced with a function that creates different negative sets, depending on the current positive pairing. This enables us to construct our Relative Contrastive Loss function in Eq. 7. This loss function iterates across all candidates, $\textbf{X}_{\textrm{i}}\in\mathcal{S}_{\textrm{cand}}$, and sets each candidate as positive, $\textbf{X}_{pos}\coloneqq\textbf{X}_{\textrm{i}}$ before calculating the negative set. All candidates with a larger distance measure than this newly defined positive are defined to be the negative set, $\mathcal{S}_{\textrm{neg}}\coloneqq f_{\textrm{neg}}(\textbf{X}_{\textrm{anc}},\textbf{X}_{\textrm{i}},\mathcal{S}_{\textrm{cand}})$, and then NT-Xent loss is calculated for this iteration.
@@
In RelCon, our pool of candidates originates from 2 sources: 1) sampling within the user, across time and 2) sampling within the batch. In (1), we choose $c$ random subsequences from the same user as the anchor sequence to be contrasted. In our experiments, $c$=20. This sampling method helps capture within-person temporal dynamics, such as how a user’s motion signals may indicate fatigue over time. Sampling within the batch in (2) allows the model to learn similarities and differences across other users. An visualization of our RelCon loss procedure can be found in Fig. 1.

### v5

@@ Abstract
We present RelCon, a novel self-supervised Relative Contrastive learning approach for training a motion foundation model from wearable accelerometry sensors [^2] Code implementation can be found at https://github.com/maxxu05/relcon. First, a learnable distance measure is trained to capture motif similarity and domain-specific semantic information such as rotation invariance. Then, the learned distance provides a measurement of semantic similarity between a pair of accelerometry time-series, which we use to train our foundation model to model relative relationships across time and across subjects. The foundation model is trained on 1 billion segments from 87,376 participants, and achieves state-of-the-art performance across multiple downstream tasks, including human activity recognition and gait metric regression. To our knowledge, we are the first to show the generalizability of a foundation model with motion data from wearables across distinct evaluation tasks.
@@ 1 Introduction
Advances in self-supervised learning (SSL) combined with the availability of large-scale datasets have resulted in a proliferation of foundation models (FMs) in computer vision (Oquab et al., 2023), language (OpenAI et al., 2023), and speech understanding (Yang et al., 2024). FMs provide powerful, general-purpose representations for a particular domain of data, and support generalization to a broad set of downstream tasks without fine-tuning. In contrast, health times-series have not yet benefited from the foundation model approach, with a few exceptions (Abbaspourazad et al., 2024; Pillai et al., 2024; McKeen et al., 2024). This is particularly unfortunate for problems in mobile health (mHealth) signal analysis, which encompasses data modalities such as IMU, PPG, and ECG, as the collection of mHealth data is both time-consuming and expensive (Rehg et al., 2017). However, recent advances in self-supervised learning for health time-series (Liu et al., 2023) have shown promise, raising the question of whether it is now feasible to train mHealth foundation models.
@@
In this paper, we demonstrate, for the first time, the feasibility of developing a foundation model for the analysis of accelerometry data across multiple tasks. Accelerometry is an important mHealth sensor used in human activity recognition (HAR) (Fu et al., 2020), physical health status assessment (Xu et al., 2022), energy expenditure estimation (Stutz et al., 2024), and gait assessment (Apple, 2021), among many other tasks. We use a novel method for self-supervised learning combined with pre-training on a large-scale, longitudinal accelerometry dataset, the Apple Heart and Movement Study (MacRae, 2021). We show that this approach yields an effective representation for accelerometry, useful for multiple downstream tasks, including HAR and gait metric estimation.
The key to our approach is a novel method for self-supervised contrastive learning, RelCon. Our approach exploits the observation that many time-series of interest are typically composed of motifs, short distinctive temporal shapes that may approximately repeat themselves over time (Mueen et al., 2009). An example is the repeating upward swing of the arm while walking, captured by a wrist-worn accelerometer. We learn a motif-based distance function and use this distance to identify positive and negative pairs sampled from within- and between-person time series for contrastive learning (Xu et al., 2024). The resulting embedding space captures the natural temporal evolution in physiological states within-person (i.e. fatigue over time) (Glass & Kaplan, 1993), and unique differences in specific physiologies between-persons (Kuncan et al., 2022; Pfeifer & Lohweg, 2018).
In this work, we introduce a novel relative contrastive loss function that models the degree of “positiveness” between the anchor and each candidate, preserving the relative relationships between instances. This helps capture diverse degrees of semantic similarity, reflecting the natural relationships found in the real-world where motions are often similar but never identical. For example, given an anchor signal of “walking,” another “walking” signal should be closest, followed by “running,” then “cycling,” and “yoga” as the most distant. Additionally, we introduce accelerometry-specific augmentations so that our distance measure can learn to compare motifs, while being invariant to the sensor’s position and orientation. Our results show that RelCon can capture subtle semantic differences between similar classes (i.e. indoor vs. outdoor cycling), as well as capture the temporally-precise information necessary for gait regression tasks. This is illustrated in Fig. 1, which shows that semantically related classes cluster more effectively in the RelCon feature representation.
The public code repository with the RelCon training set-up, architecture, and evaluation can be found here: https://github.com/maxxu05/relcon. Our key contributions are:
  1. 1.
    We introduce RelCon, a novel contrastive learning methodology that models relative differences between instances via a learnable motif-based distance measure. This aims to model the varying degrees of semantic similarity found in real-world motion data.
  2. 2.
    We utilize RelCon to pre-train a foundation model on 1b accelerometry segments across 87k participants from the Apple Heart and Movement Study. This large-scale, longitudinal dataset, sourced from real-world wearable data, enables modeling of human motion dynamics across diverse contexts, supporting robust applications in health and activity monitoring.
  3. 3.
@@
    Our RelCon-trained motion foundation model achieves state-of-the-art performance against 11 models across 6 unique downstream datasets with 4 different tasks, ranging from classification of 16 in-the-wild activities to regression of temporally-precise gait metrics. We are the first to show the generalizability of a motion foundation model across distinct, diverse evaluations.
@@ 2 Related Work
Figure 2: SOTA Accelerometry SSL Methods. Each sequence color represents a different user’s time-series. RelCon draws candidates from both within- and between-user and ranks them by their relative similarity via a learnable distance function. Then, it iteratively applies a contrastive loss, selecting one candidate as positive while assigning the more distant as negative. This helps prevent false positives/negatives because the full relative ranking is captured. Prior approaches define a single positive/negative set, risking semantic errors if positive/negatives are misdefined. AugPred and SimCLR are resistant to false positives because they construct positive pairs via semantics-preserving augmentations, but REBAR does not have a semantically-constrained pair construction.
@@
Motion time-series FM: We believe we are the first to train a foundation model for accelerometry data. We define foundation models to be representations that are pre-trained on broad-scale data and are capable of solving multiple diverse downstream tasks without fine-tuning (Bommasani et al., 2021) (e.g., each task uses frozen weights with supervised training of a light-weight prediction head). The closest related work is Yuan et al. (2024), which uses 700k person-days of data from UK Biobank to train an accelerometry representation. While their model is evaluated on multiple datasets, all of the tasks are HAR, making it unclear whether the learned representation captures broader motion information. In contrast, we test our FM on multiple gait analysis tasks as well as HAR, demonstrating the effectiveness of our representation for both regression and classification tasks. Narayanswamy et al. (2025) also focuses on HAR but uses minute-level hand-crafted features instead of the raw 100 Hz signals necessary for temporally-precise motion tasks, such as fall detection (Sucerquia et al., 2018) or gait assessment (Alobaidi et al., 2022). Other notable FMs focus on general time-series forecasting (Das et al., 2023; Ansari et al., 2024) or do not model accelerometry motion data (Goswami et al., 2024; Abbaspourazad et al., 2024; Saha et al., 2025; Song et al., 2024).
Motion time-series SSL: Our RelCon architecture introduces a novel contrastive learning approach for time-series data. It is most closely-related to REBAR’s  (Xu et al., 2024) motif-based positive pair generation approach. However, REBAR is not adapted to accelerometry and lacks invariance to well-understood accelerometry-semantic-preserving transformations. Additionally, REBAR employs a “hard” contrastive loss, leading to sensitivity to false positives and negatives due to rigidly designating one positive and all others as negatives. RelCon overcomes these limitations by learning a motif-based distance with accelerometry-specific augmentations and introducing a relative loss function to mitigate false labeling effects. Most prior motion SSL approaches focus exclusively on HAR (Logacjov, 2024; Haresamudram et al., 2024; Straczkiewicz et al., 2021). Haresamudram et al. (2022) benchmarked a number of SSL approaches on accelerometer data, testing generalizability across datasets and sensor positions. The best methods focus on either utilizing augmentations in a SimCLR-fashion (Tang et al., 2020) or utilizing them as prediction targets (Saeed et al., 2019). Comparison of these two models to REBAR can be seen in Fig. 2. Models for other motion-based topics – including walking speed (Shrestha & Won, 2018; Soltani et al., 2021), gait (Slemenšek et al., 2023; Brand et al., 2024), and health monitoring (Takallou et al., 2024) have focused on single-task supervised learning. For alternative biosignals, such as PPG (Pillai et al., 2024), ECG (Jeong et al., 2023; Diamant et al., 2022), and EEG (Pradeepkumar et al., 2025), there have been recent works that develop domain-specific SSL approaches.
@@
Soft Label Learning Approaches: A feature of our RelCon approach is the use of soft labels to obtain more fine-grained characterizations of similarity. Most prior self-supervised contrastive methods for time-series have adopted the hard label approach  (Yue et al., 2022; Tonekaboni et al., 2021; Zhang et al., 2022), including the previous REBAR method (Xu et al., 2024). One exception is Lee et al. (2023), which proposed a soft contrastive approach for classification and anomaly detection. However, unlike RelCon, they utilize non-learnable distance measures, such as dynamic time warping (Müller, 2007), that cannot be trained to capture domain-specific semantic features.
@@ 3 Methodology
The RelCon methodology has two key components. The first is in its innovations to learn a distance measure that captures semantic accelerometry information, described in Section 3.1. The second is a novel relative contrastive loss that encodes relative order relationships, described in Section 3.2.
@@ 3.1 Learnable Distance Measure
where $\smash{\textbf{X}\in\mathbb{R}^{T\times D}}$ and $\smash{\textbf{x},\mu,\sigma\in\mathbb{R}^{D}}$ with $T$ as the time length and $D$=3 for our 3-axis accelerometry signals. The distance between an anchor sequence and a candidate sequence, $\smash{d(\textbf{X}_{\textrm{anc}},\textbf{X}_{\textrm{cand}})}$, is defined as the reconstruction accuracy to generate the anchor from the candidate. The distance measure is strictly dependent on the motif similarities between the anchor and candidate that are captured in the dilated convolutions in $\smash{f_{\{q/k\}}}$ (Xu et al., 2024). See code for further details.
@@
During training, we define $\smash{\textbf{X}_{\textrm{cand}}\coloneqq Aug(\textbf{X}_{\textrm{anc}})}$, making the candidate an augmented version of the original sequence, using augmentation procedure from Accel SimCLR (Tang et al., 2020). This prevents an exact-match solution and helps the model learn to reconstruct the original anchor from semantically similar but altered candidates, such as recognizing running signals despite variations like an upside-down wearable or noise from a loose-fitting device.
@@
    Use reversible instance normalization (Kim et al., 2021) to normalize an anchor based upon the candidate, preserving relative anchor-candidate magnitude information.
The anchor sequence and final reconstruction output are normalized (in Eq. 4) and unnormalized (in Eq. 2) using candidate sequence statistics. When an anchor and candidate have drastically different magnitudes, it is more likely they represent different fundamental motions and should have increased reconstruction error. Reversible normalization helps preserve this effect.
@@ 3.2 Relative Contrastive Loss: All pairs are positive …
Normalized temperature cross entropy loss (NT-Xent) is standard in contrastive learning (Eq. 5). It learns a class discriminative embedding space by pulling the anchor and positive instances closer together in the embedding space and pushing all negatives away from the anchor, as in:
@@
where $\smash{\textrm{sim}(\textbf{X}_{\textrm{anc}},\textbf{X}_{\textrm{pos}})}=\langle E(\textbf{X}_{\textrm{anc}}),E(\textbf{X}_{\textrm{pos}})\rangle$ and $E(\cdot)$ is some encoder function. Traditionally, NT-Xent uses strict, “hard” labels that define absolute positive and negative sets. However, we would like to use the learned distance measure to capture relative ordering among the candidates in the loss. That is, if $\smash{d(\textbf{X}_{\textrm{anc}},\textbf{X}_{i})>d(\textbf{X}_{\textrm{anc}},\textbf{X}_{j})}$, the loss learns to preserve that relative ordering in the $E(\cdot)$’s embedding space. To achieve this, we redefine $\smash{\mathcal{S}_{\textrm{neg}}\coloneqq f_{\textrm{neg}}(\textbf{X}_{\textrm{anc}},\textbf{X}_{\textrm{i}},\mathcal{S}_{\textrm{cand}})}$ as a function in Eq. 6, allowing the set of negatives to dynamically change based on the current positive pair. This enables the construction of our Relative Contrastive Loss function in Eq. 7:
@@
This loss iteratively calculates NT-Xent over each candidate, $\smash{\textbf{X}_{\textrm{i}}\in\mathcal{S}_{\textrm{cand}}}$, iteratively setting each candidate as the positive, $\smash{\textbf{X}_{pos}\coloneqq\textbf{X}_{\textrm{i}}}$. The negative set for the given positive is then defined as all candidates with a greater distance measure than the given positive $\smash{\{\textbf{X}\in\mathcal{S}:d(\textbf{X}_{\textrm{anc}},\textbf{X})>d(\textbf{X}_{\textrm{anc}},\textbf{X}_{\textrm{pos}})\}}$. Candidates with larger distances from the anchor appear in more negative sets, reinforcing the relative ranking. In this way, the RelCon loss function captures hierarchical structure of similarities in the embedding space based on their relative relationships with the anchor.
In order to capture diverse relationships via RelCon, our pool of candidates originates from two sources: (1) sampling within the user, across time and (2) sampling between users, within the batch. In (1), we choose $c$ random subsequences from the same user as the anchor sequence. This helps capture within-person temporal dynamics in the encoder, such as how a user’s motion signals may indicate fatigue over time. In (2), sampling between-users allows the model to learn physiological similarities and differences across other users. Our RelCon procedure is visualized in Fig. 2.